On a trip of 800 km, Charlie drove at two different speeds. In one province he drove at 100 km/h and in another at 110 km/h. If the entire trip took seven and one-half hours, the distance travelled at the 110 km/h speed is?
Answer by stanbon(48551)

On a trip of 800 km, Charlie drove at two different speeds. In one province he drove at 100 km/h and in another at 110 km/h. If the entire trip took seven and one-half hours, the distance travelled at the 110 km/h speed is?
----
1st: Province DATA:
distance = x km ; rate = 100 km/h ; time = d/r = x/100 hrs
----------------------
2nd: Provice DATA:
distance = (800-x) km ; rate = 110 km/h ; time = d/r = (800-x)/110 hrs
---------------------
Equation:
time + time + time
x/100 + (800-x)/110 = 7.5 hrs
110x + 100(800-x) = 100*110*7.5
110x + 100*800 - 100x = 100*825
10x = 100*25
x = 250 km
----
Find distance traveled at 110 km:
800-x = 800-250 = 550 km
======================================
Cheers,
Stan H.